OpenStudy (anonymous):

Help please! Which equation defines the graph of y = x^2 after it is shifted vertically 3 units down and horizontally 5 units left? y = (x - 5)^2 - 3 y = (x + 5)^2 - 3 y = (x - 3)^2 - 5 y = (x + 3)^2 - 5

OpenStudy (anonymous):

Nope. /:

OpenStudy (anonymous):

That makes sense, but would would that make my answer out of my answer choices? (5,3) isn't an option. y = (x - 5)^2 - 3 y = (x + 5)^2 - 3 y = (x - 3)^2 - 5 y = (x + 3)^2 - 5 These are the choices.

OpenStudy (anonymous):

y=x^2 means (y-0)=(x-0)^2 where (0,0) is the vertex of your parabola after shifting your graph it is (-5,-3) so equation is [y-(-3)]=[x-(-5)]^2

OpenStudy (anonymous):

y=(x+5)^2-3

OpenStudy (anonymous):

second one